### v1

# Gravitational Effects of a Small Primordial Black Hole Passing Through the Human Body

Abstract
The gravitational effects of a primordial black hole (PBH) passing through the human body are examined, with the goal of determining the minimum mass necessary to produce significant injury or death. Two effects are examined: the damage caused by a shock wave propagating outward from the black hole trajectory, and the dissociation of brain cells from tidal forces produced by the black hole on its passage through the human body. It is found that the former is the dominant effect, with a cutoff mass for serious injury or death of approximately $M_{PBH}>1.4\times 10^{17}{\rm g}$. The number density of primordial black holes with a mass above this cutoff is far too small to produce any observable effects on the human population.
I Introduction
The detection of gravitational radiation from merging black holes LIGO as well as the direct imaging of M87* M87* and Sgr A* SgrA* has led to a significant resurgence of interest in black hole physics. Among the ideas that have been reexplored is the possibility that primordial black holes (PBH) could account for some or all of the dark matter in the Universe Carr(2020) . Many constraints can be placed on the mass of black holes that can serve as dark matter; the main open mass window lies in the asteroid mass range Carr(2021) ; Gorton(2024)
$10^{17}{\rm g}<M_{PBH}<10^{22}{\rm g}.$ (1)
Here Hawking radiation provides the lower bound: any PBHs with a mass below $10^{15}$ g would have evaporated by the present day, while limits on the PBH decay products push this lower bound up to about $10^{17}$ g. The upper bound in Eq. (1) is due to limits on black hole microlensing events, but other limits rule out black hole dark matter at higher masses. Any black holes lying in the mass range given by Eq. (1) must be primordial, as they cannot be produced through normal stellar evolution.
If primordial black holes exist and are sufficiently abundant, they could produce a variety of interesting effects. Several authors have explored the consequences of a PBH impact on the Earth Khriplovich(2008) ; Luo(2012) ; Rahvar(2021) , and in Ref. Luo(2012) it was argued that a PBH as small as $10^{15}$ g would produce an easily-detectable seismic disturbance. Dai and Stojkovic Dai have suggested that PBHs could accumulate inside of planets and asteroids and hollow out their cores. They also pointed out that the passage of a PBH through material on the earth would leave a detectable microscopic tunnel and suggested further that such a passage through the human body would not be fatal. It is this latter question that we address here.
Clearly, a sufficiently large PBH would have a significant impact on the human body; the goal of this present work is to determine the minimum black hole size needed to produce significant injury, and to see whether the observed lack of such injuries can place any bounds on PBH properties. Our discussion is partially motivated by earlier work on macroscopic dark matter (MACROs) Sidhu(2020) . In that paper, it was shown that the impact of MACRO dark matter on the human body would be sufficiently destructive that the non-observance of such impacts allows useful limits to be placed on the MACRO mass and cross section. While it is well known that a microscopic black hole would have a deleterious effect on the human body, there are no precise calculations. Most previous discussions of this topic are in the popular literature Loeb(2021) ; in fact, the first reference to the effect of a PBH on the human body appears to be in a work of fiction Niven(1974) .
In the next section, the biological effects of a PBH passing through the human body are discussed, including both the shock wave produced by such a passage and the tidal forces that would tear apart human cells. The minimum mass needed to produce serious injury or death from each of these effects is derived. In Sec. III we examine whether such effects can place useful limits on PBH properties. While the minimum mass needed to produce serious injury or death lies within the interesting asteroid mass range, the allowed PBH abundance in this mass range is too small to produce any observable effects on the human population.
II Primordial black holes and the human body
[…]
Consider first the shock wave generated by a PBH passing through the human body. Khripolovich et al. Khriplovich(2008) estimated the total energy that would be deposited by a PBH in the form of a shock wave as it passed through the Earth. Their estimate for the total energy deposition is Khriplovich(2008)
$E=4\pi(GM_{PBH}/v_{PBH})^{2}\rho L\ln(c_{s}/\sqrt{4\pi G\rho}a).$ (2)
Here $M_{PBH}$ and $v_{PBH}$ are the PBH mass and velocity, $\rho$ is the density of the target, $L$ is the path length of the PBH through the target, $c_{s}$ is the sound speed in the target, and $a$ is the typical intermolecular distance in the target. We can take the human body to be primarily water, with $\rho=1$ g/cm3, $c_{s}=1.5\times 10^{5}$ cm/s, and $a=3\times 10^{-8}$ cm. Then the logarithmic factor on the right-hand side of Eq. (2) is 36.
How much energy would need to be deposited in the human body to produce significant injury or death? The shock wave produced in this process would be similar to that produced by a bullet passing through the human body, which is a well-studied process. In the latter case, the tissue damage scales roughly as the deposited energy Courtney(2012) ; Baum(2022) . Following Ref. Sidhu(2020) , we will take the threshold for serious energy to be approximately 100 J, which is the typical muzzle energy of a .22 caliber rifle Rhee(2016) . Using this value for $E$ in Eq. (2), along with a typical path length through the human body of $L=$ 10 cm and a PBH velocity on the order of the dark matter velocity dispersion ($v_{PBH}\sim 200$ km/s), we find that the lower bound on $M_{PBH}$ to produce significant injury in this fashion is
$M_{PBH}>1.4\times 10^{17}{\rm g}.$ (3)
Now consider a second mechanism for tissue destruction. As the PBH passes through the human body, it will exert strong tidal forces in its vicinity. These will produce a tensile force on nearby human cells, and a sufficiently strong force would tear the cells apart. The cells most sensitive to this dissociation are likely to be those in the human brain. A tensile force of $10-100$ nN for a few microseconds would be sufficient to pull apart human brain cells Miller(2002) ; Hutson(2024) .
Recall that the tidal force on a cell with mass $m_{cell}$ and typical size $r_{cell}$ at a distance $d$ from the PBH is given by
$F=\frac{2GM_{PBH}m_{cell}r_{cell}}{d^{3}}.$ (4)
Note that the brain can be surprising resilient to damage over very small regions. Hence, we will err on the side of conservatism and require that the entire brain be subject to a tidal force of $10-100$ nN. This will insure sufficient destruction of brain tissue (basically, the entire brain) to produce a detectable effect. Furthermore, a PBH velocity of  200 km/s is equivalent to 20 cm/$\mu$s, so that it will take $\sim\mu$s for the PBH to traverse the entire brain, ensuring that these tidal forces will be applied to the entire brain for at least a few microseconds. We will take $r_{cell}\sim 10~{}\mu$m, and $m_{cell}\sim\rho r_{cell}^{3}$ with $\rho$ again taken to be the density of water. Then with $d\sim 10$ cm, we find that $F\sim 10-100$ nN corresponds to a lower bound on the PBH mass of
$M_{PBH}>7\times 10^{18}{\rm g}-7\times 10^{19}{\rm g}.$ (5)
Even given the crudeness of our approximations, it is clear that the effect of the gravitationally-induced shock wave on the human body is larger than the effect of the tidal forces. Thus, Eq. (3) gives the smallest PBH mass that will produce significant damage to the human body from gravitational forces.
III Conclusions
Intriguingly, the smallest PBH mass that can lead to significant human injury (Eq. 3) lies near the lower bound on PBH dark matter in Eq. (1), raising the possibility that that the absence of such events could provide an additional constraint on the PBH abundance, as in the case of MACHO dark matter Sidhu(2020) . Unfortunately, it is easy to see that the effect of PBH collisions with the human population will always be negligible. Following Ref. Sidhu(2020) , we estimate the rate of PBH-human interactions as
$\Gamma=\frac{\rho_{PBH}N_{human}A_{human}v_{PBH}}{M_{PBH}},$ (6)
where $N_{human}$ is the total human population ($N_{human}=8\times 10^{9}$) and $A_{human}$ is the typical cross-sectional area of a human ($A_{human}\sim 1~{}{\rm m}^{2}$). Taking $\rho_{PBH}$ to account for all of the dark matter ($\rho_{PBH}=2\times 10^{-30}$ g/cm3), and using the smallest mass in Eq. (3) that allows for serious injury, we obtain $\Gamma\sim 10^{-18}$ events per year, which is utterly negligible.
This result is not suprising, as there are two major differences between PBHs and the model considered in Ref. Sidhu(2020) . The first is that the MACHO model allows for much smaller dark matter masses and correspondingly larger number densities, resulting in a higher interaction rate with the human population. The second is that MACHOs are strongly interacting, while PBHs interact gravitationally, and gravity is, of course, notoriously weak. Perhaps the more interesting result is just how weak the limit in Eq. (3) turns out to be. A black hole with $M_{PBH}\sim 10^{16}$ g corresponds to the mass of a small asteroid, but such a black hole could pass right through the human body and cause only negligible gravitational effects.
Acknowledgements.
I thank A. Loeb, M.S. Hutson, and J.R. Scherrer for helpful discussions.